# 3D-free meets 3D priors: Novel View Synthesis from a Single Image with Pretrained Diffusion Guidance

### v4

@@ 5.1 Qualitative analysis
Our model is able to generate distinct viewpoints in the camera control to ensure consistency across generated views. Details are mentioned in the Zero123++ Shi et al. (2023a). We present qualitative representative results and comparisons with Zero123++ Shi et al. (2023a), HawkI Kothandaraman et al. (2023b), and StableZero123 at camera angles of $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ in Figures 5 and 6. Our method demonstrates superior scene reconstruction from all viewpoints compared to previous works. Specifically, results on HawkI-Syn in Figure 5 show that StableZero123 is largely ineffective. HawkI fails to capture the correct camera elevation in all cases except for the house image. While Zero123++ handles both elevation and azimuth, it struggles with background and detailed features. For instance, the pyramid in the first row lacks shadow information; the waterfall image in the second row appears unnatural; and the house in the last row blurs detailed features. Conversely, our model accurately reflects shadow characteristics in the pyramid, and reconstructs the details and background of the waterfall and house examples from various viewpoints.
Similar observations are made for HawkI-Real results shown in Figure 6. StableZero123 is ineffective. Zero123++ fails to capture background or detailed information. For example, when tasked with camera control for an image of the Eiffel Tower, Zero123++ focuses solely on the Eiffel Tower, ignoring surrounding details. The original HawkI model, while producing aerial views, fails in angle conversion tasks. In contrast, our model accurately performs angle conversion tasks at $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$, including the Seine River in the background for the Eiffel Tower image, showcasing its superiority. Camera control tasks for the HawkI-Real dataset, including images like the Hawaii beach and a cat, further demonstrate our model’s excellence compared to other models. The key benefits of our model over 3D-based NVS methods such as Zero123++ and 3D-free methods such as HawkI arises by merging the strengths of 3D-based techniques into a 3D-free optimization process, effectively combining the best features of both.
5.2 Quantitative Evaluation
@@
Following prior work Shi et al. (2023a); Kothandaraman et al. (2023b); Liu et al. (2023a), we evaluate our method using six metrics - (i) LPIPS Zhang et al. (2018): Quantifies the perceptual similarity between the generated and input images, with lower values indicating better performance. (ii) CLIP-Score Radford et al. (2021) and View-CLIP Score: Measure text-based alignment of the generated images. The CLIP score assesses alignment with both content and the $(\alpha_{elev},\alpha_{azi})$ viewpoint, while the View-CLIP Score focuses specifically on the viewpoint. Higher values are preferred. (iii) DINO Caron et al. (2021), SSCD Pizzi et al. (2022): DINO evaluates the semantic consistency of the generated images by comparing high-level feature embeddings extracted from a self-supervised vision transformer. DINO is trained not to ignore differences between subjects of the same class. Higher values indicate better preservation of semantic content across different views of the same scene. SSCD measures structural similarity between the generated images and their reference counterparts using learned feature representations. SSCD focuses on capturing fine-grained structural and contextual consistency. Higher values are preferred for better alignment with ground-truth structures. (iv) CLIP-I Ruiz et al. (2023): CLIP-I measures the cosine similarity between the embeddings of multi-view images and the input image within the CLIP space. (v) PSNR (Peak Signal-to-Noise Ratio) and SSIM Wang et al. (2004) (Structural Similarity Index): PSNR Quantifies the pixel-wise fidelity of the generated images relative to the reference images. PSNR is calculated as the logarithmic ratio of the maximum possible pixel value to the mean squared error between the two images. Higher values indicate better pixel-level accuracy. SSIM assesses perceptual similarity by comparing luminance, contrast, and structural information between the generated and reference images. SSIM is designed to measure structural consistency, with higher values reflecting closer perceptual alignment.
Similar to the quantitative comparison performed by Zero123++, we use 10% of the overall data from the HawkI-Syn and HawkI-Real datasets as the validation set to compute the quantitative metrics. Table 1 and Table 5 shows that our model significantly outperforms the state-of-the-art across these evaluation metrics, reinforcing how our models stands out by incorporating the robust features of 3D-based NVS methods into a 3D-free optimization strategy, thereby capitalizing on the benefits of both approaches.
@@
5.3 Ablation analysis: viewpoint regularization loss
To demonstrate the effectiveness of our approach in achieving camera control, we present results for scenes such as a Hawaiian beach and a waterfall. In both instances, the guidance images from Zero123++ fail to provide accurate direction to the model. For the Hawaiian beach scene, the output generated with the regularization term exhibits a more consistent style compared to the output produced without it. Despite the inaccuracies in the Zero123++ guidance images, the regularization term facilitates more reliable camera control than the results generated without it.
Similarly, in the waterfall scene, the regularization term enhances the consistency of the generated rock textures surrounding the waterfall. Without the regularization term, these textures are inconsistently represented; however, with it, the style is maintained more faithfully. Once again, Zero123++ does not provide accurate guidance in this case, underscoring the significant contribution of the regularization loss to improved control and visual coherence in the generated images. Detailed results from our ablation study are presented in Figure 7. Furthermore, the application of the regularization loss demonstrates performance improvements in quantitative evaluations, as shown in Table 2.
6 Conclusions, Limitations and Future Work
Appendix A Appendix
A.1 Additional Results
A.2 Computation Time
Model Memory Consumption Computation Time
Zero123++ 10.18 GB / 40.0 GB (9,715 MiB) 20 sec
Stable Zero123 39.3 GB / 40.0 GB (37,479 MiB) 1,278 sec
ZeroNVS 33.48 GB / 40.0 GB (31,929 MiB) 7,500 sec
Model Step Memory Consumption Computation Time
HawkI Optimization 7.20 GB / 40.0 GB (6,875 MiB) 395 sec
Inference 8.43 GB / 40.0 GB (8,045 MiB) 6 sec (each image)
Total 8.43 GB / 40.0 GB (8,045 MiB) 401 sec
Ours (w/o regloss) Zero123++ 10.18 GB / 40.0 GB (9,715 MiB) 20 sec
Optimization 7.21 GB / 40.0 GB (6,879 MiB) 372 sec
Inference 8.43 GB / 40.0 GB (8,049 MiB) 6 sec (each image)
Total 10.18 GB / 40.0 GB (9,715 MiB) 398 sec
Ours (w/ regloss) Zero123++ 10.18 GB / 40.0 GB (9,715 MiB) 20 sec
Optimization 7.21 GB / 40.0 GB (6,885 MiB) 367 sec
Inference 8.43 GB / 40.0 GB (8,045 MiB) 6 sec (each image)
Total 10.18 GB / 40.0 GB (9,715 MiB) 393 sec
A.3 Quantitative Evaluation Results

### v5

@@ Introduction
Our model is able to generate distinct viewpoints in the camera control to ensure consistency across generated views. Details are mentioned in the Zero123++ (Shi et al. 2023a). We present qualitative representative results and comparisons with Zero123++ (Shi et al. 2023a), HawkI (Kothandaraman et al. 2023b), and StableZero123 at camera angles of $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ in Figures 5 and 6. Our method demonstrates superior scene reconstruction from all viewpoints compared to previous works. Specifically, results on HawkI-Syn in Figure 5 show that StableZero123 is largely ineffective. HawkI fails to capture the correct camera elevation in all cases except for the house image. While Zero123++ handles both elevation and azimuth, it struggles with background and detailed features. For instance, the pyramid in the first row lacks shadow information; the waterfall image in the second row appears unnatural; and the house in the last row blurs detailed features. Conversely, our model accurately reflects shadow characteristics in the pyramid, and reconstructs the details and background of the waterfall and house examples from various viewpoints. Similar observations are made for HawkI-Real results shown in Figure 6. StableZero123 is ineffective. Zero123++ fails to capture background or detailed information. For example, when tasked with camera control for an image of the Eiffel Tower, Zero123++ focuses solely on the Eiffel Tower, ignoring surrounding details. The original HawkI model, while producing aerial views, fails in angle conversion tasks. In contrast, our model accurately performs angle conversion tasks at $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$, including the Seine River in the background for the Eiffel Tower image, showcasing its superiority. Camera control tasks for the HawkI-Real dataset, including images like the Hawaii beach and a cat, further demonstrate our model’s excellence compared to other models. The key benefits of our model over 3D-based NVS methods such as Zero123++ and 3D-free methods such as HawkI arises by merging the strengths of 3D-based techniques into a 3D-free optimization process, effectively combining the best features of both.
Quantitative Evaluation
@@
Following prior work (Shi et al. 2023a; Kothandaraman et al. 2023b; Liu et al. 2023a), we evaluate our method using six metrics - (i) LPIPS (Zhang et al. 2018): Quantifies the perceptual similarity between the generated and input images, with lower values indicating better performance. (ii) CLIP-Score (Radford et al. 2021): Measure text-based alignment of the generated images. The CLIP score assesses alignment with both content and the $(\alpha_{elev},\alpha_{azi})$ viewpoint. Higher values are preferred. (iii) DINO (Caron et al. 2021), SSCD (Pizzi et al. 2022): DINO evaluates the semantic consistency of the generated images by comparing high-level feature embeddings extracted from a self-supervised vision transformer. DINO is trained not to ignore differences between subjects of the same class. Higher values indicate better preservation of semantic content across different views of the same scene. SSCD measures structural similarity between the generated images and their reference counterparts using learned feature representations. SSCD focuses on capturing fine-grained structural and contextual consistency. Higher values are preferred for better alignment with ground-truth structures. (iv) CLIP-I (Ruiz et al. 2023): CLIP-I measures the cosine similarity between the embeddings of multi-view images and the input image within the CLIP space. (v) PSNR (Peak Signal-to-Noise Ratio) and SSIM (Wang et al. 2004) (Structural Similarity Index): PSNR Quantifies the pixel-wise fidelity of the generated images relative to the reference images. PSNR is calculated as the logarithmic ratio of the maximum possible pixel value to the mean squared error between the two images. Higher values indicate better pixel-level accuracy. SSIM assesses perceptual similarity by comparing luminance, contrast, and structural information between the generated and reference images. SSIM is designed to measure structural consistency, with higher values reflecting closer perceptual alignment. Similar to the quantitative comparison performed by Zero123++, we use 10% of the overall data from the HawkI-Syn and HawkI-Real datasets as the validation set to compute the quantitative metrics. Table 1 and Table 3 shows that our model significantly outperforms the state-of-the-art across these evaluation metrics, reinforcing how our models stands out by incorporating the robust features of 3D-based NVS methods into a 3D-free optimization strategy, thereby capitalizing on the benefits of both approaches.
@@
Ablation analysis: viewpoint regularization loss
To demonstrate the effectiveness of our approach in achieving camera control, we present results for scenes such as a Hawaiian beach and a waterfall. In both instances, the guidance images from Zero123++ fail to provide accurate direction to the model. For the Hawaiian beach scene, the output generated with the regularization term exhibits a more consistent style compared to the output produced without it. Despite the inaccuracies in the Zero123++ guidance images, the regularization term facilitates more reliable camera control than the results generated without it. Similarly, in the waterfall scene, the regularization term enhances the consistency of the generated rock textures surrounding the waterfall. Without the regularization term, these textures are inconsistently represented; however, with it, the style is maintained more faithfully. Once again, Zero123++ does not provide accurate guidance in this case, underscoring the significant contribution of the regularization loss to improved control and visual coherence in the generated images. Detailed results from our ablation study are presented in Figure 7. Furthermore, the application of the regularization loss demonstrates performance improvements in quantitative evaluations, as shown in Table 2.
Conclusions, Limitations and Future Work
Appendix A Appendix
Computation Time
Model Memory Consumption Computation Time
Zero123++ 10.18 GB / 40.0 GB (9,715 MiB) 20 sec
Stable Zero123 39.3 GB / 40.0 GB (37,479 MiB) 1,278 sec
ZeroNVS 33.48 GB / 40.0 GB (31,929 MiB) 7,500 sec
Model Step Memory Consumption Computation Time
HawkI Optimization 7.20 GB / 40.0 GB (6,875 MiB) 395 sec
Inference 8.43 GB / 40.0 GB (8,045 MiB) 6 sec (each image)
Total 8.43 GB / 40.0 GB (8,045 MiB) 401 sec
Ours (w/o regloss) Zero123++ 10.18 GB / 40.0 GB (9,715 MiB) 20 sec
Optimization 7.21 GB / 40.0 GB (6,879 MiB) 372 sec
Inference 8.43 GB / 40.0 GB (8,049 MiB) 6 sec (each image)
Total 10.18 GB / 40.0 GB (9,715 MiB) 398 sec
Ours (w/ regloss) Zero123++ 10.18 GB / 40.0 GB (9,715 MiB) 20 sec
Optimization 7.21 GB / 40.0 GB (6,885 MiB) 367 sec
Inference 8.43 GB / 40.0 GB (8,045 MiB) 6 sec (each image)
Total 10.18 GB / 40.0 GB (9,715 MiB) 393 sec
Additional Results